Ie a car is slsina docn a a rate of x-orad/s2, hou many revolutions, around a arcular track of radios 200m, would it do if it & seaso Slats nrth a speed of ug,-Trad aud it takes him 20s to sop. Assume eo dalafeulate the sinal angle et b. Calculate the Length of the ciraula track do? C. How many Turns wao it alble to ow many, lurns wa。iT able to
Answer #1

Solution:

A) The final angle,

Using the concep of angular kinmatics,

?2f - ?20 = 2???

02 - 42 = 2*(-6)*?f

?f = 4/3 = 1.333 radian.

B) The lenght of the ciruclar track is,

C = Circumfarence of circle= 2?r = 2?*200 = 1256.64 m

C) Distace covered by the car is,

? = 1.333 radians = S/r

S = 1.333*200 = 266.667 m

There fore number o turns is,

n = S/C = 0.212 turn.

I hope you understood the problem, If yes rate me!! or else comment for a better solution.